• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ramesh

    Log entry time 08:09:45 on March14,2005

    Entry number 141453

    keyword=Shift Summary


    HRS angles: left 19.5 right 40
    HRS momentum settings: left 3.762 right 1.45
    BigBite and Neutron Array angles : both -99

    00:00 Shift started with no beam

    Took controlled access (during the no beam time) to solve the problem of HV communication with Epics, Brad swapped ethernet cable to see if it was due to the bad cable. It did not work.

    01:10 Beam back

    01:25 Finished Harp scan , see 141419 .

    All production runs on the Target: tilted carbon at 40uA

    01:27-02:08 Run 2641 (200k)
    02:08-02:52 Run 2642 (200k)
    02:52-03:28 Run 2643 (200k)
    03:30-04:09 Run 2644 (200k)
    04:10-04:25 Run 2645 (junk) DAQ problem.

    04:25 DAQ stopped working (deadtime was 100 in 41uA also). ND-scaler display also showed no reading (no cosmic either), ND/BB-HV was ok (seen in the port server screen). After rebooting all VME's of the DAQ everything worked fine.

    04:39- Run 2646 (junk) ROC6 problem

    05:00 Roc6 failed (disconnected status), SRC Fastbus and SRC trigger supervisor reset, rebooted ROC6,did "startcoda", worked fine.

    05:43-05:05 Run 2647 (200k)
    05:51-06:40 Run 2648 (200k)
    06:40-07:18 Run 2649 (200k)
    07:23- Run 2650 passed to day shift.